# First-order and Berezinskii-Kosterlitz-Thouless phase transitions in two-dimensional generalized XY models

### v1

@@ Abstract
The aim of this paper is to illustrate that generalized two-dimensional XY models (proposed by Romano and Zagrebnov) may also support a first-order phase transition. Two approaches are employed to accurately determine the critical parameter $q$ at which such a transition takes place. Furthermore, we show that the model is characterized by three distinct regions concerning both first-order and Berezinskii-Kosterlitz-Thouless phase transitions. Finally, the underlying mechanisms governing such transitions are presented, along with an estimation of the critical temperatures.
@@
Figure 1: (a) Energy for spin in function of temperature for different values of q, and $L=96$. We can observe a discontinuity being formed with the increase of q.Specific heat for q=3 (b), and the scaling of max specific heat with $L^{2}$ (c), showing comportament of a first order transition.
@@
Two-dimensional magnetic models characterized by short-range interaction and continuous symmetry fail to exhibit long-range order at a finite temperature [1]. Despite the absence of long-range order due to the destructive influence of spin waves, even at low temperatures, a quasi-long-range order with a power-law decaying correlation function is evident. At higher temperatures, correlations decay exponentially, leading to a disordered phase. The transition from quasi-long-range order to a disordered phase constitutes a topological phase transition within the Berezinskii–Kosterlitz–Thouless (BKT) universality class [2, 3]. This transition is characterized by the unbinding of topological objects, know as vortices, at a critical temperature $T_{BKT}$.
@@
Among the simplest models supporting topological excitations in the spin field and a BKT transition are those described by the Hamiltonian $H=-J\sum_{<ij>}(S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j})$, where $J$ is a coupling constant, the summation is over nearest-neighbors sites in a square lattice, and $S^{\alpha}_{i}$ represents the spin at site $i$. From this Hamiltonian one can extract the Planar Rotator Model (PRM) by enforcing the constraint $S^{2}_{x}+S^{2}_{y}=1$, which confines the spins to the plane. On the other hand, this same Hamiltonian leads to the XY-model, simply by allowing spins to have three components and the constraint $S^{2}_{x}+S^{2}_{y}+S^{2}_{z}=1$. Both models can be parameterized by using scalar fields. For the PRM, the azimuthal angle $\phi$ is enough, yielding the Hamiltonian $H=-j\sum_{<ij>}\cos(\phi_{i}-\phi_{j})$. Meanwhile, for the XY model, both azimuthal and polar angles are employed, resulting in the Hamiltonian $H=-j\sum_{<ij>}\sin\theta_{i}\sin\theta_{j}\cos(\phi_{i}-\phi_{j})$.
@@
Extensive studies on the static critical behavior of such models have established consensus regarding phase transitions, critical temperature and exponents [4] [5]. The nature of the phase transitions was very investigated in the 1970’s and 1980’s, such that different generalizations and extensions of the these two models have become also object of studies.
@@
Domany et al. [6] introduced a generalization of the PRM model utilizing the Villain form $V(\phi)=1-\cos^{p^{2}}(\phi^{\prime}/2)$, with $\phi^{\prime}=\phi_{i}-\phi_{j}$. As $p$ increases, a narrow potential well with width $\pi/p$ emerges, resembling the $n$-states Potts model with $p$ proportional to $n$. With Monte Carlo simulations, Domany et al. [6] revealed that an increase in $p$ leads to a discontinuity in energy and a two-peak structure in the energy histogram, suggesting a first-order transition for large $p$. Some studies challenged this result, particularly those based on renormalization group analysis [7, 8]. However, the possibility of first-order induced disorder in the planar model was demonstrated [9], and later, rigorous proof by van Enter and Shlosman confirmed that $SO(n)$ invariant $n$-vector models with sufficiently deep and narrow minima can undergo a first-order phase transition[10].
@@
A generalized XY-model was proposed by Romano and Zagrebnov[11], whose Hamiltonian is given by,
$H^{Gen}_{XY}=$ $-J\sum_{<i,j>}\big{[}1-(S^{z}_{i}S^{z}_{i})^{2}-(S^{z}_{j}S^{z}_{j})^{2}$ (1)
@@
Writing this with the help of the two scalar fields $\theta$ and $\phi$, one gets,
$H^{Gen}_{XY}=-J\sum_{<i,j>}(\sin\theta_{i}sin\theta_{j})^{q}\cos\phi_{i}\cos\phi_{j}.$ (2)
Here, $q\in$ N is the generalization parameter, and for $q=1$, the XY model is recovered. In their investigations[11], Romano and Zagrebnov used rigorous inequalities for all values of $q$, considering systems in two and three dimensions. For the two-dimensional (2D) case, they showed that, for arbitrary $q$, the model has orientational disorder at all finite temperatures, undergoing a transition to a low temperature phase with slow decay of correlations and infinite susceptibility. So the class of universality of the transition is not necessarily the same as that of the proper BKT transition. Mól et al. [12] extensively studied the vortex-like solutions within the generalized XY model, employing both the continuous Hamiltonian and the Self Consistent Harmonic Approximation (SCHA) [13]. They showed that only static planar-vortex configurations are stable and supported for any $q\geq 1$, and that the critical temperature of phase transitions decreases with the parameter $q$. In addition, Monte Carlo simulations [14] suggest a first-order phase transition in the generalized XY model for lower values of $q$, but the result was not conclusive. Although the first order transition was initially contested, a rigorous proof by van Enter et al. [15] confirmed that the 2D generalized XY model proposed by Romano and Zagrabenov indeed exhibits a first-order phase transition, leaving open the value of the critical parameter $q_{c}$ at which this transition begins to appear.
Our motivations for the present work are threefold: firstly, we would like to confirm the existence of a first-order phase transition for Hamiltonian (1); the second motivation is to obtain the generalized critical parameter $q_{c}$ where the system exhibits a first-order transition; finally, our task concentrates in providing a physical interpretation for this first-order phase transition. For these purposes, we have used the hybrid Monte Carlo algorithm, not much different from that used by Mól et al. [14], which will be explained briefly in the next section. To enhance readability, most thermodynamic results and techniques for determining transition temperatures are presented in appendices and only the main results are shown in the usual text. We begin by presenting Monte Carlo simulation results and then, guided by these calculations and the solution of van Enter et al. [15], we analytically estimate the the critical temperature of the first-order phase transition.
@@ II Monte Carlo Method
Figure 2: (a) show how the gap varies with the size of the system, comparing the two sizes of L=80,96.(b) show how the transition varies with the temperature for q=3.(c) Lee-Kosterlitz criterion for q=3,5,6, where the expected comportament of a first order transition is observed for the three cases.
We have employed a Hybrid Monte Carlo approach to generate configurations for calculating thermal averages. A Monte Carlo Step (MCS) involves a sequence of operations, starting with a Wolff cluster algorithm for in-plane components [16], followed by modified Metropolis single-spin updates for the three components and $N$ over-relaxation steps. In the modified Metropolis algorithm [17], a randomly selected spin undergoes a small increment in random components, followed by renormalization to maintain unit length. The acceptance is determined by the standard Metropolis algorithm, with the spin increment length chosen to achieve an acceptance rate between 30% and 60%. This modification proved effective in mitigating critical slowdown at low temperatures. Various acceptance rates were tested, with little impact on the results. Over-relaxation involved reflecting a randomly selected spin across the effective magnetic field due to its neighbors while preserving the spin length. This algorithm effectively alters configurations while maintaining energy conservation. Mean values were obtained using 10,000 Monte Carlo passes to warm up the samples, and 300,000 configurations were individually divided into each bin for analysis. Testing higher numbers of configurations and bin sizes showed minimal impact on the results.
@@ III.1 Monte Carlo Results
Initially, we will examine the impact of the generalized parameter $q$ on the energy per spin ($e$) and the specific heat ($c_{v}$). For the simulations involving the usual XY-model ($q=1$), our results align with expectations [18]. Both energy and specific heat exhibit minor finite-size effects as we approach the critical temperature; the estimated errors escalate due to rising fluctuations and critical slowing down. A distinct peak emerges at a temperature exceeding the critical temperature of the BKT transition. The determination of this critical temperature involves various methods, detailed in table 1 and further explicated in the appendix. With the increasing of the parameter $q$, the specific heat peak moves toward lower temperatures (still above the critical temperature observed for a BKT transition) and becomes narrow and higher as the system size $L$ increases. The energy begins to exhibit a discontinuity (Fig 1a) at the same temperature where the maximum specific heat ($c_{v}^{max}$) occurs. These two characteristics hint at the possibility of a first-order phase transition. However, to definitively characterize a first-order transition from simulations, it is imperative to observe that the transition persists in the thermodynamic limit ($L\rightarrow\infty$). Regarding the specific heat [19], the maximum value exhibits a proportional relationship with the volume of the system dimension, expressed as $c_{v}^{max}\propto L^{d}$. Therefore, plotting $c_{v}^{max}$ versus $L^{2}$ for various values of $q$ can reveal if they present a linear dependence. As Fig.1c shows, for $q=1$, no finite-size effect is observed. However, for $q=3$ and higher, the specific heat peak continues to increase with the size of the system, but obtaining the precise maximum value becomes increasingly challenging due to fluctuations.
@@
A more effective approach for discerning a first-order phase transition involves computing the histogram of the energy distribution [20]. This is achieved through extended runs within the temperature region surrounding the energy discontinuity for different lattice sizes. In a first-order transition, the theory predicts a double-peak structure [21], each peak representing a distinct state of the system: one ordered, the other disordered. The transition temperature is defined when two peaks with equal areas emerge in the histogram, indicating equiprobability between the two states. For any $q<3$, no double-peak structure was identified for any lattice size $L$, leading to the conclusion that the only one phase transition for the system is of the BKT type. However, for $q=3$, a subtle double-peak histogram is observed, particularly for $L=64$, with a temperature transition difference of less than 1% compared to the energy discontinuity and the specific heat peak temperature. As $L$ increases, the double-peak structure becomes more apparent as shown in 2, because the free energy barrier between the two states becomes higher. The double-peak structure becomes more pronounced with increasing $L$ (see Fig. 2), reflecting a growing free energy barrier between the two states and, again, the transition temperature differs by less than 1% from the energy discontinuity and specific heat peak temperatures.
@@
To definitively confirm the first-order transition, we employ the Lee-Kosterlitz criterion [22]. According to this reasoning, determining a first-order transition involves calculating the free energy barrier, denoted as $\triangle F(L)$, for various sizes of the system $L$. The Lee-Kosterlitz criterion implies that $\triangle F(L)$ should be independent of $L$, while for a first-order phase transition, it is expected to be an increasing function of $L$. So, it is crucial to highlight the dependence on the size of the lattice. For $q=3$, the double-peak structure shows up faintly for $L=64$. For higher values of $q$, the double-peak structure manifests much earlier; for instance, in $q=4$ one can observe it for $L=32$. This inherent size dependence poses challenges in applying the Lee-Kosterlitz criterion to small values of $q$ and the same problem was observed by Lee-Kosterlitz [22] in their own paper for small values of potts model. They attributed this issue to the weakness of the transition. However, for $q=3,5,6$, we observe the anticipated behavior indicative of a first-order phase transition, as depicted in Fig. 2c. Despite computational limitations prevent the acquisition of the Lee-Kosterlitz criterion, the observed increasing in the free energy barrier supports the assertion that the critical parameter is $q=3$. In Table 1, we show the values of the transition temperature utilizing the Binder Fourth Cumulant (see Complementary material). This method reveals another characteristic of a first-order phase transition: a jump to a negative value precedes the attainment of the value $U_{L}=0.5$ for an ordered system [23].
@@
In the range from $q=3$ to $q=5$, the critical temperature obtained through finite-size scaling of the susceptibility and the Helicity ( see complementary material) consistently reveals transition temperatures higher than the actual transition temperature. However, for $q=6$ and higher, the critical temperature and the transition temperature obtained from the Binder Fourth Cumulant and the susceptibility converge. Consequently, the determination of the critical temperature becomes challenging based on the finite-size scaling of the susceptibility and Binder Fourth Cumulant. This observation potentially suggests the absence of the BKT transition for this generalized parameter.
@@
The only one method that has shown possible to determine a critical temperature is through the finite-size scaling (FSS) of the helicity modulus. This is due to the system not only exhibiting vortices but also becoming progressively populated with the increase of the parameter $q$, as illustrated in 3.
@@
Figure 3: Vortex Density in function of temperature for different values of q, and $L=96$. We can observe a discontinuity being formed with the increase of q.
@@
The vortex density exhibits a discontinuity with increasing $q$ at the same temperature as the energy and where the specific heat presents a peak, indicating a correlation between these events. To illustrate this correlation more effectively, we calculate the vortex density difference. Essentially, at the system transitions between states, the vortex density difference jumps from 00 to a non-zero value. To better demonstrate this point, we have taken the vortex density difference as an order parameter, that is, when the system change from one state for another, the vortex density jumps from 00 to a specified value. This fact allowed us to define a way to obtain the critical generalization parameter of the model, simply by calculating the mean vortex density difference for different $q$ at the region of temperature where supposedly the discontinuity should appear. The first value in which the mean vortex density difference is appreciable is found to be $q=3$, agreeing with the result obtained by the histogram and Binder Cumulant.
@@
Figure 4: Two distinct configurations were observed for $q=6$ at the system’s transition temperature, $T_{c}=0.6212$, with a lattice size of $L=96$. The red points denote vortices, while black points represent antivortices. The state labeled as ”Disordered” with higher energy, exhibits a significantly larger population of vortices compared to the ”Ordered” state with lower energy.
III.2 Gibbs Measures and First Order Transitions

### v2

@@ Abstract
Besides the Berezinskii-Kosterlitz-Thouless phase transition, the two-dimensional generalized XY model, identified by a generalization parameter $q$ (as proposed by Romano and Zagrebnov), can also support a first-order phase transition, starting from a critical value $q_{c}$. However, the value of $q_{c}$ at which this transition takes place is unknown. In this paper, we take two approaches to determine the critical parameter $q_{c}$ accurately. Furthermore, we show that the model is characterized by three distinct regions concerning both first-order and Berezinskii-Kosterlitz-Thouless phase transitions. Finally, the underlying mechanism governing such transitions is presented, along with an estimation of the critical temperatures.
@@
Figure 1: (a) Energy per spin as a function of temperature for different values of $q$ and $L=96$. As $q$ increases, we can observe a discontinuity being formed. Specific heat for $q=3$ (b), and the scaling of max specific heat with $L^{2}$ (c), showing the behavior of a first-order phase transition.
@@
Two-dimensional magnetic models characterized by short-range interaction and continuous symmetry fail to exhibit magnetic long-range order at a finite temperature [1]. Despite the absence of magnetic long-range order due to the destructive influence of spin waves, even at low temperatures, a quasi-long-range order with a power-law decaying correlation function may occur. At higher temperatures, correlations always decay exponentially, leading to a disordered phase. The transition from quasi-long-range order to a disordered phase constitutes a topological phase transition within the Berezinskii–Kosterlitz–Thouless (BKT) universality class [2, 3]. This transition is characterized by the unbinding of topological objects, known as vortices, at a critical temperature $T_{BKT}$.
@@
Among the simplest models supporting topological excitations in the spin field and a BKT transition are those described by the Hamiltonian $H=-J\sum_{<ij>}(S^{x}_{i}S^{x}_{j}+S^{y}_{i}S^{y}_{j})$, where $J$ is a coupling constant, the summation is over nearest-neighbors sites in a square lattice, and $S^{\alpha}_{i}$ represents the spin at site $i$. From this Hamiltonian one can extract the Planar Rotator Model (PRM) by enforcing the constraint $S^{2}_{x}+S^{2}_{y}=1$, which confines the spins to the plane. On the other hand, this same Hamiltonian leads to the XY model, simply by allowing spins to have three components and the constraint $S^{2}_{x}+S^{2}_{y}+S^{2}_{z}=1$. Both models can be parameterized by using scalar fields. For the PRM, the azimuthal angle $\phi$ is enough, yielding the Hamiltonian $H=-J\sum_{<ij>}\cos(\phi_{i}-\phi_{j})$. Meanwhile, for the XY model, both azimuthal and polar angles are employed, resulting in the Hamiltonian $H=-J\sum_{<ij>}\sin\theta_{i}\sin\theta_{j}\cos(\phi_{i}-\phi_{j})$.
@@
Extensive studies on the static critical behavior of such models have established consensus regarding phase transitions, critical temperature, and exponents [4] [5]. The nature of the phase transitions was intensively investigated in the 1970s and 1980s, such that different generalizations and extensions of these two models have become objects of study.
@@
Domany et al. [6] introduced a generalization of the PRM model utilizing the Villain model $V(\phi)=1-\cos^{p^{2}}(\phi^{\prime}/2)$, with $\phi^{\prime}=\phi_{i}-\phi_{j}$. As p increases, a narrow potential well with width $\pi/p$ emerges, resembling the n-states Potts model with p proportional to n. With Monte Carlo simulations, Domany et al. [6] revealed that an increase in p leads to a discontinuity in energy and a two-peak structure in the energy histogram, suggesting a first-order transition for large p. Some studies challenged this, particularly those based on renormalization group analysis [7, 8]. However, the possibility of first-order transition between disordered phases induced in the planar model was demonstrated [9], and later rigorous proof by Van Enter and Shlosman confirmed that SO(n) invariant $n$-vector models with sufficiently deep and narrow minima can undergo a first-order phase transition[10].
@@
A generalized XY model was proposed by Romano and Zagrebnov[11], whose Hamiltonian is given by,
$H^{Gen}_{XY}=$ $-J\sum_{<i,j>}\big{[}1-(S^{z}_{i})^{2}-(S^{z}_{j})^{2}$ (1)
@@
Writing this with the help of the two scalar fields $\theta$ and $\phi$, one gets
$H^{Gen}_{XY}=-J\sum_{<i,j>}(\sin\theta_{i}\sin\theta_{j})^{q}\cos(\phi_{i}-\phi_{j}).$ (2)
Here, $q$ is the generalization parameter and the XY model is recovered for $q=1$. Although $q$ does not need to be an integer number for the Hamiltonian model to be well-defined, in this work, we will consider only integer numbers for $q$ as defined by the original articles. In their paper[11] they used rigorous inequalities for all values of $q$ for systems in two and three dimensions. For the two-dimensional (2D) case, they showed that, for arbitrary $q$, the model has orientational disorder at all finite temperatures, and undergoes a transition to a low-temperature phase with slow decay of correlations and infinite susceptibility. In thermodynamic terms, the class of universality of the transition may or may not correspond to that of the proper BKT transition. Mól et al. [12] extensively studied the vortex-like solutions within the generalized XY model, employing both the continuous Hamiltonian and the Self Consistent Harmonic Approximation (SCHA) [13]. They showed that only static planar-vortex configurations are stable and supported for any $q\geq 1$ and that the critical temperature of phase transitions decreases with the parameter $q$. In addition, Monte Carlo simulations [14] suggest a first-order phase transition in the generalized XY model for sufficiently high values of $q$, but the result was not conclusive. Although the first-order transition was initially contested, a rigorous proof by van Enter et al. [15] confirmed that the 2D generalized XY model proposed by Romano and Zagrebnov indeed exhibits a first-order phase transition, leaving open the value of the critical parameter $q_{c}$ at which this transition begins to appear.
Above, we have seen part of the effort to understand the generalized XY model. However, there are still several open questions regarding the two different phase transitions and their possible coexistence in the system. Here, our main motivation is to establish the correct values of the generalized critical parameter $q_{c}$, above which a first-order phase transition takes place. It is a fundamental step for further investigations related to the model. For that, we have used the hybrid Monte Carlo algorithm, not much different from that used by Mól et al. [14], which will be explained briefly in the next section. To enhance readability, most thermodynamic results and techniques for determining transition temperatures are presented in appendices, with only the main results in the usual text. We begin by presenting Monte Carlo simulation results and then, guided by these calculations and the solution of van Enter et al. [15], we analytically estimate the critical temperature of the first-order phase transition.
@@ II Monte Carlo Method
Figure 2: (a) shows how the gap varies with the size of the system, comparing the two sizes of $L=80$,$96$.(b) the two peak-structure is easily observed for $q=6$ and $L=96$ at the transition temperature. (c) Lee-Kosterlitz criterion for $q=3$,$5$,$6$, where the expected behavior of a first-order transition is observed for the three cases.
We have employed a Hybrid Monte Carlo approach to generate configurations for calculating thermal averages. A Monte Carlo Step (MCS) involves a sequence of operations, starting with a Wolff cluster algorithm for in-plane components [16], followed by modified Metropolis single-spin updates for the three components and N over-relaxation steps. In the modified Metropolis algorithm [17], a randomly selected spin underwent a small increment in random components, followed by renormalization to maintain unit length. The acceptance was determined by the standard Metropolis algorithm, with the spin increment length chosen to achieve an acceptance rate between 30% and 60%. This modification proved effective in mitigating critical slowdown at low temperatures. Several acceptance rates were tested, with little impact on the results. Over-relaxation involved reflecting a randomly selected spin across the effective magnetic field due to its neighbors while preserving the spin length. This algorithm effectively altered configurations while maintaining energy conservation. Mean values were obtained using 10,000 Monte Carlo passes to warm up the samples, and 300,000 configurations were individually divided into each bin for analysis. Testing higher numbers of configurations and bin sizes has shown minimal impact on the results.
@@ III.1 Monte Carlo Results
Initially, we will examine the impact of the generalization parameter $q$ on the energy for spin ($e$) and the specific heat ($c_{v}$). For the simulations involving the usual XY model ($q=1$), our results align with expectations [18]. Both energy and specific heat exhibit minor finite-size effects as we approach the critical temperature; the estimated errors escalate due to rising fluctuations and critical slowing down. A distinct peak emerges at a temperature exceeding the critical temperature of the BKT transition. The determination of this critical temperature involves various methods, detailed in table 1 and further explained in the appendix. With the increasing of the parameter $q$, the specific heat peak moves toward lower temperatures (still above the critical temperature observed for a BKT transition) and becomes narrow and higher as the system size $L$ increases. The energy begins to exhibit a discontinuity (fig. 1a) at the same temperature where the maximum specific heat ($c_{v}^{max}$) occurs. These two characteristics hint at the possibility of a first-order phase transition. However, to definitively characterize a first-order transition from simulations, it is imperative to observe that the transition persists in the thermodynamic limit ($L\rightarrow\infty$). Regarding the specific heat [19], the maximum value exhibits a proportional relationship with the volume of the system dimension, expressed as $c_{v}^{max}\propto L^{d}$. Therefore, plotting $c_{v}^{max}$ versus $L^{2}$ for various values of $q$ can reveal if they present a linear dependence. As shown in Fig 1c, for $q=1$, no finite-size effect is observed. On the other hand, for q=3 and higher, the specific heat peak increases with the size of the system. However, getting the precise maximum value of the peak becomes increasingly challenging due to fluctuations.
@@
A more effective approach for discerning a first-order phase transition involves computing the histogram of the energy distribution [20]. This is achieved through extended runs within the temperature region surrounding the energy discontinuity for different lattice sizes. In a first-order transition, the theory predicts a double-peak structure [21], each peak representing a distinct state of the system: one ordered, the other disordered. The transition temperature is defined when two peaks with equal height emerge in the histogram, indicating equiprobability between the two states. For any $q<3$, no double-peak structure was identified for any lattice size $L$, leading to the conclusion that the only phase transition for the system is of the BKT type. However, for $q=3$, a subtle double-peak histogram was observed, particularly for $L=64$, with a transition temperature difference of less than 1% compared to the energy discontinuity and the specific heat peak temperature. As $L$ increases, the double-peak structure becomes more apparent as shown in Fig. 2, because the free energy barrier between the two states becomes higher. The double-peak structure becomes more pronounced with increasing $L$ (see Fig. 2), reflecting a growing free energy barrier between the two states and, again, the transition temperature differs by less than 1% from the energy discontinuity and specific heat peak temperatures.
@@
To confirm the first-order transition, we will employ the Lee-Kosterlitz criterion [22]. According to this criterion, determining a first-order transition involves calculating the free energy barrier $\triangle F(L)$ across different system sizes $L$. So, the criterion implies that $\triangle F(L)$ should be independent of $L$, and at a first-order phase transition, it is expected to be an increasing function of $L$. For $q=3$, the double-peak structure shows up faintly for $L=64$. For higher values of $q$, the double-peak structure manifests itself much earlier, for example, in $q=4$ we already can observe it for $L=32$. This inherent size dependence poses challenges in applying the Lee-Kosterlitz criterion to small values of $q$. A similar problem was observed by Lee and Kosterlitz [22] in their paper concerning the Potts model with small values of the generalization parameter. They attributed this issue to the weakness of the transition. However, for $q=3,5,6$, we observe the anticipated behavior indicative of a first-order phase transition, as depicted in Fig. 2c. Despite computational limitations preventing the application of the Lee-Kosterlitz criterion, the observed increase in the free energy barrier supports the assertion that the critical parameter is $q=3$. In Table 1, we show the values of the transition temperature utilizing the Binder Fourth Cumulant (see Appendix B). This method reveals another characteristic of a first-order transition: a jump to a negative value precedes the attainment of the value $U_{L}=0.5$ for an ordered system [23].
@@
In the range from $q=3$ to $q=5$, the critical temperature obtained through finite-size scaling of the susceptibility and the Helicity (see Appendix C and Appendix D respectively) consistently reveals transition temperatures higher than the critical temperature obtained. However, for $q=6$ and higher, the critical temperature and the transition temperature obtained from the Binder Fourth Cumulant and the susceptibility converge. Consequently, the determination of the critical temperature becomes challenging based on the finite-size scaling of the susceptibility and Binder Fourth Cumulant. This observation potentially suggests the absence of the BKT transition for this generalized parameter.
@@
To determine the critical temperature, the only method that was possible to determine a critical temperature was through the finite-size scaling (FSS) of the helicity modulus. This is because the system exhibits vortices and the population of these topological objects increases rapidly as the generalization parameter $q$ increases, as illustrated in 3.
@@
Figure 3: Vortex Density as a function of temperature for different q values, and $L=96$. We can observe a discontinuity being formed with the increase of q.
@@
The vortex density exhibits a discontinuity for sufficiently large $q$ and such discontinuity occurs at the same temperature as the energy and where the specific heat presents a peak, indicating a correlation between these events. To illustrate this correlation more effectively, we calculated the vortex density difference as an order parameter. Essentially, when the system transitions between states, the vortex density difference jumps from 0 to a non-zero value. Such behavior allows us to define a way to get the critical generalization parameter of the model, simply by calculating the mean vortex density difference for different q at the region of temperature where supposedly the discontinuity should appear. The first value where the mean vortex density difference is appreciable is given by $q=3$, which agrees with the result obtained by the histogram and Binder Cumulant.
@@
Figure 4: Two distinct configurations were observed for $q=6$ at the system’s transition temperature, $T_{c}=0.6212$, with a lattice size of $L=96$. The red points denote vortices, while the black points represent antivortices. The state labeled ”Disordered” with higher energy, exhibits a significantly larger population of vortices compared to the ”Ordered” state with lower energy.
III.2 Gibbs Measures and First-Order Transitions